Group Selection Page HELP Controls Group ITDM Vacuum WP78 SPB SQS XO ToBeMigrated2XWiki ToBeMigrated2GitlabPages ToBeMigratedByGroup Migrated2Zulip Archived2PDF not2MigrateOrArchive
SCS SXP Migrated2Zulip Photon Commissioning Team XRO DOC HED Migrated2Zulip WP74 MID PSPO_XO XO_Operation SQS Migrated2Zulip FXE Migrated2Zulip DET Sample Environment SPB Migrated2Zulip
OPERATION-2025 OPERATION-2024 OPERATION-2023 OPERATION-2022 OPERATION-2021 OPERATION-2020 OPERATION-2019 SASE1-2018 SASE1-2017 Experiment data
  OPERATION-2024 logbook, Contact: F.Wolff-Fabris, A.Galler, H.Sinn  Not logged in ELOG logo
Message ID: 151     Entry time: 22 Aug 2024, 23:15
Author: Raphael de Wijn 
Group: SPB 
Subject: Shift summary 

X-ray delivery

12.40 - 12.82 keV, ~1.1 mJ

Optical laser delivery

ns laser for jet visualization
ns laser for pumping ready but not used

Achievements/Observations

  • Focus IRD
  • ns time and spatial overlap
  • Collected lyzo data
  • Collected dark rhodopsin data
  • New opaque foil installed on cube X-ray exit side

Issues

  • Roadrunner motor issues solved after manual intervention and computer restart
ELOG V3.1.4-7c3fd00